### v3

@@ 4.3 Evaluation Procedure
[…]
@@ 4.4 Results
You are given a data file at ‘assets/grades.txt’ that contains multiple entries representing individuals and their respective grades. Fix the incorrect regex between ### FIX CODE BELOW and ### FIX CODE ABOVE to return a list of just those individuals who received a grade of ‘B’.
The first few lines of the data file are illustrated below:
@@
Given two strings S1 and S2 as input, the task is to merge them alternatively, i.e., the first character of S1 then the first character of S2 and so on till the strings end.
Expected Time Complexity: $O(|S1|+|S2|)$.
@@
1#User function Template for python3 2class Solution: 3  def merge(self, S1, S2): 4    l1=len(S1) 5    l2=len(S2) 6    if l1>l2: 7      res="" 8      dif=l1-l2 9      x=S1[:l2] 10      y=S2 11      d=[(i,j) for i,j in zip(x,y)] 12      for i in d: 13        for j in i: 14          res+=j 15      return res+S1[-dif:] 16    elif l1<l2: 17      res="" 18      dif=l2-l1 19      x=S2[:l1] 20      y=S1 21      d=[(i,j)for i,j in zip(x,y)] 22      for i in d: 23        for j in i: 24          res+=j 25      return res+S2[-dif:] 26    else: 27      res="" 28      x=S1[:l2] 29      y=S2 30      d=[(i,j)for i,j in zip(x,y)] 31      for i in d: 32        for j in i: 33          res+=j 34      return res 35 36#{ Driver Code Starts 37#Initial Template for Python 3 38if __name__ == ’__main__’: 39    t = int(input()) 40    for _ in range(t): 41        S1,S2 = map(str,input().strip().split()) 42        ob = Solution() 43        print(ob.merge(S1, S2)) 44# } Driver Code Ends (b) Student’s buggy program 1#User function Template for python3 2class Solution: 3  def merge(self, S1, S2): 4    l1=len(S1) 5    l2=len(S2) 6    if l1>l2: 7      res="" 8      dif=l1-l2 9      x=S1[:l2] 10      y=S2 11      d=[(i,j) for i,j in zip(x,y)] 12      for i in d: 13        for j in i: 14          res+=j 15      return res+S1[l2:] 16    elif l1<l2: 17      res="" 18      dif=l2-l1 19      x=S1 20      y=S2[l1:] 21      d=[(i,j)for i,j in zip(x,y)] 22      for i in d: 23        for j in i: 24          res+=j 25      return res+S2[l1:] 26    else: 27      res="" 28      x=S1 29      y=S2 30      d=[(i,j)for i,j in zip(x,y)] 31      for i in d: 32        for j in i: 33          res+=j 34      return res 35 36#{ Driver Code Starts 37#Initial Template for Python 3 38if __name__ == ’__main__’: 39    t = int(input()) 40    for _ in range(t): 41        S1,S2 = map(str,input().strip().split()) 42        ob = Solution() 43        print(ob.merge(S1, S2)) 44# } Driver Code Ends (c) Fixed program For Input: Qh    eyNFX
@@
The bug in the program is in the slicing of the strings when one string is longer than the other. In the original program, when S1 is longer than S2, the remaining characters of S1 are not correctly added to the result. The same issue occurs when S2 is longer than S1. The fix is to slice the longer string from the length of the shorter string to the end, instead of slicing from the end. (e) Detailed explanation Consider how you are slicing the longer string when one string is longer than the other. (f) Single-sentence hint Refer to caption (g) Validation

### v4

@@ 4.3 Evaluation Procedure
Default with trials $\textit{k}{}=3$, GPT-3.5 student model, use $\mathcal{X}$, and $94.7~{}(0.0)$ 76.076.0\ \ 76.076.0 $97.6~{}(2.4)$ 87.587.5\ \ 87.587.5 $95.5~{}(4.5)$ 73.373.3\ \ 73.373.3
@@ 4.4 Results
You are given a data file at ‘assets/grades.txt’ that contains multiple entries representing individuals and their respective grades. Fix the incorrect regex between ### FIX CODE BELOW and ### FIX CODE ABOVE to return a list of just those individuals who received a grade of ‘B’. The first few lines of the data file are illustrated below:
@@
(b) Student’s buggy program
[…]
(c) Fixed program
Your Code’s
@@
(e) Detailed explanation
Consider how the regular expression groups are defined and how they affect the output format.
(f) Single-sentence hint
(g) Validation
@@
Given two strings S1 and S2 as input, the task is to merge them alternatively, i.e., the first character of S1 then the first character of S2 and so on till the strings end. Expected Time Complexity: $O(|S1|+|S2|)$.
@@
1#User function Template for python3 2class Solution: 3 def merge(self, S1, S2): 4 l1=len(S1) 5 l2=len(S2) 6 if l1>l2: 7 res="" 8 dif=l1-l2 9 x=S1[:l2] 10 y=S2 11 d=[(i,j) for i,j in zip(x,y)] 12 for i in d: 13 for j in i: 14 res+=j 15 return res+S1[-dif:] 16 elif l1<l2: 17 res="" 18 dif=l2-l1 19 x=S2[:l1] 20 y=S1 21 d=[(i,j)for i,j in zip(x,y)] 22 for i in d: 23 for j in i: 24 res+=j 25 return res+S2[-dif:] 26 else: 27 res="" 28 x=S1[:l2] 29 y=S2 30 d=[(i,j)for i,j in zip(x,y)] 31 for i in d: 32 for j in i: 33 res+=j 34 return res 35 36#{ Driver Code Starts 37#Initial Template for Python 3 38if __name__ == ’__main__’: 39 t = int(input()) 40 for _ in range(t): 41 S1,S2 = map(str,input().strip().split()) 42 ob = Solution() 43 print(ob.merge(S1, S2)) 44# } Driver Code Ends
(b) Student’s buggy program
1#User function Template for python3 2class Solution: 3 def merge(self, S1, S2): 4 l1=len(S1) 5 l2=len(S2) 6 if l1>l2: 7 res="" 8 dif=l1-l2 9 x=S1[:l2] 10 y=S2 11 d=[(i,j) for i,j in zip(x,y)] 12 for i in d: 13 for j in i: 14 res+=j 15 return res+S1[l2:] 16 elif l1<l2: 17 res="" 18 dif=l2-l1 19 x=S1 20 y=S2[l1:] 21 d=[(i,j)for i,j in zip(x,y)] 22 for i in d: 23 for j in i: 24 res+=j 25 return res+S2[l1:] 26 else: 27 res="" 28 x=S1 29 y=S2 30 d=[(i,j)for i,j in zip(x,y)] 31 for i in d: 32 for j in i: 33 res+=j 34 return res 35 36#{ Driver Code Starts 37#Initial Template for Python 3 38if __name__ == ’__main__’: 39 t = int(input()) 40 for _ in range(t): 41 S1,S2 = map(str,input().strip().split()) 42 ob = Solution() 43 print(ob.merge(S1, S2)) 44# } Driver Code Ends
(c) Fixed program
For Input: Qh    eyNFX
@@
The bug in the program is in the slicing of the strings when one string is longer than the other. In the original program, when S1 is longer than S2, the remaining characters of S1 are not correctly added to the result. The same issue occurs when S2 is longer than S1. The fix is to slice the longer string from the length of the shorter string to the end, instead of slicing from the end.
(e) Detailed explanation
Consider how you are slicing the longer string when one string is longer than the other.
(f) Single-sentence hint
(g) Validation